# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) has outpaced traditional evaluation methodologies. It presents novel challenges, such as measuring human-like psychological constructs, navigating beyond static and task-specific benchmarks, and establishing human-centered evaluation. These challenges intersect with Psychometrics, the science of quantifying the intangible aspects of human psychology, such as personality, values, and intelligence. This survey introduces and synthesizes an emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs. We systematically explore the role of Psychometrics in shaping benchmarking principles, broadening evaluation scopes, refining methodologies, validating results, and advancing LLM capabilities. This paper integrates diverse perspectives to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field. Ultimately, we aim to provide actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit. A curated repository of LLM psychometric resources is available at https://github.com/valuebyte-ai/Awesome-LLM-Psychometrics.
@@ 1 Introduction
"Whatever exists at all exists in some amount. To know it thoroughly involves knowing its quantity as well as its quality." [Thorndike, 1962]
@@
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
Traditional AI evaluation has relied on curating task-specific datasets, annotating ground-truth labels with human input, running models on these datasets, and assessing performance using predefined metrics. However, LLMs have triggered an evaluation crisis, as their versatile capabilities and human-like behaviors exceed what traditional benchmarks can measure. Novel challenges include evaluating psychological constructs like personality, values, and cognitive biases that conventional assessment cannot capture. Rapid LLM development and training data contamination have made static benchmarks obsolete. Additionally, the prompt- and context-sensitive nature of LLMs compromises the robustness and validity of existing evaluation frameworks. With increasing human-LLM interaction, human-centered evaluation approaches are becoming essential. Moreover, as LLMs integrate into agentic and multimodal systems, evaluation methodologies must expand in scope and complexity.
@@
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology, including knowledge, skills, character, values [Pasquali, 2009]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014]. The convergence of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
@@
We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (such as personality, values, morality, and attitudes) and cognitive constructs (including heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities). Research in LLM psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs. Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor. By measuring and elucidating psychological constructs in LLMs, LLM psychometrics further informs strategies for their targeted enhancement.
@@
The field of LLM psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic fields. Despite this diversity, there is a lack of cohesion among researchers from different communities, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs. Consequently, there is a pressing need for a comprehensive survey to synthesize these efforts and facilitate a more integrated understanding of the field.
@@
This paper aims to bridge the gap by providing the first systematic review of LLM psychometrics, encompassing evaluation, validation, and enhancement. Fig. 1 illustrates the structure of the review. § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. In § 3, we delineate the definition, scope, and taxonomy of LLM psychometrics, which establishes the structure of the remainder of the review. The core measurement framework is detailed in § 5, § 6, and § 7. § 5 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In § 6, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 7, which examines the psychometric validation of evaluation results. Beyond evaluation, § 8 introduces strategies for enhancing LLMs through psychometric insights. Finally, § 9 discusses emerging trends, challenges, and future directions in the psychometric evaluation of LLMs, while § 10 concludes the paper.
Related Surveys.
We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024]. Several related surveys focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025]. Hagendorff [2023], Hagendorff et al. [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic survey of LLM psychometrics.
@@ 2.3 Psychometric Evaluation of AI Before the Era of LLMs
3 LLM Psychometrics: Definition, Scope, and Taxonomy
We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
LLM psychometrics research adapts, advances, and innovates upon psychometric methodologies to rigorously evaluate LLMs. This evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). In many cases, these psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve).
The remainder of this paper begins by thoroughly contrasting psychometrics and traditional AI benchmarking (§ 4.1), identifying key differences such as core goals, philosophical foundations, measurement methods, and result analysis (Table 1). We then review how recent LLM psychometrics research adopts and extends psychometric principles to underpin and reinvent LLM benchmarks (§ 4.2).
We proceed by examining the principal constructs assessed within LLM psychometrics (what to measure; § 5), including both personality (§ 5.1) and cognitive constructs (§ 5.2). Personality constructs are categorized into personality traits, values, morality, and attitudes and opinions, while cognitive constructs comprise heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities. For each construct, we synthesize the underlying theoretical frameworks and measurement instruments, critically compare alternative approaches, and distill the main empirical findings.
We further scrutinize the psychometric evaluation methodologies employed for LLMs (how to measure; § 6). The methodological framework is systematically analyzed in terms of its key components: test format (§ 6.1), data and task sources (§ 6.2), prompting strategies (§ 6.3), and model output and scoring (§ 6.4). In addition, we discuss the role of inference parameters in shaping evaluation results (§ 6.5).
We next address the psychometric validation of measurement results, examining both reliability and validity (how well do we measure; § 7). Reliability is assessed using metrics such as internal consistency, parallel forms reliability, and inter-rater agreement (§ 7.1). Validity is considered from content, construct, criterion, and ecological perspectives (§ 7.2). Persistent concerns include prompt sensitivity, data contamination, response biases, and divergence from human psychological constructs. To mitigate these challenges, recent research advocates for standardized, procedural tests, including multiple task versions, prompt perturbation, and transparent reporting practices, thereby promoting rigorous, reproducible, and meaningful psychometric evaluation of LLMs (§ 7.3).
The psychometric insights further inform the development and refinement of LLMs (how to improve; § 8). Current research predominantly focuses on enhancing LLMs’ performance along three dimensions: trait manipulation (§ 8.1), safety and alignment (§ 8.2), and cognitive enhancement (§ 8.3). Finally, we discuss the current trends, challenges, and future directions of LLM psychometrics (§ 9).
@@ 5 Psychometrics for Measuring Psychological Constructs
HEXACO Wang et al. [2025], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
@@
ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al.
@@ 5.1 Measuring Personality Constructs
LLMs exhibit personality constructs that are not explicitly programmed or trained towards. The emergent constructs critically shape LLM behavior [Ye et al., 2025a, Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024]. Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
Table 2 summarizes the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, Li et al. [2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
@@ 5.1.1 Personality Traits
"Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life." [APA Dictionary of Psychology, n.d.]
Personality traits define an individual’s patterns of thinking, feeling, and behaving [Larsen et al., 2005]. In the context of LLMs, Zhang et al. [2024a], Wang et al. [2025] have explored how personality traits relate to model safety, bias, and toxicity. Other works emphasize that chatbot personality highly determines user experience [Jiang et al., 2023, Serapio-García et al., 2023, Huang et al., 2024c, Klinkert et al., 2024].
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences. The Big Five model [Goldberg, 2013] identifies five core dimensions of personality: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Building upon the Big Five, the HEXACO model [Ashton and Lee, 2007] introduces an additional dimension, Honesty-Humility, which encompasses traits such as sincerity, fairness, and modesty. Another influential framework, the Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], categorizes individuals into 16 distinct personality types based on four dichotomous dimensions: Extraversion vs. Introversion, Sensing vs. Intuition, Thinking vs. Feeling, and Judging vs. Perceiving. In contrast to these models, the Dark Triad [Paulhus and Williams, 2002] focuses on three socially aversive personality traits—Machiavellianism, Narcissism, and Psychopathy—highlighting the darker aspects of human behavior.
@@
When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits [Caron and Srivastava, 2022, Huang et al., 2023d, a, Karra et al., 2022, La Cava and Tagarelli, 2024, Serapio-García et al., 2023]; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits [Miotto et al., 2022, Wang et al., 2025]; MBTI assessment [Myers, 1985] for MBTI types [Cui et al., 2023, Huang et al., 2023c, La Cava and Tagarelli, 2024, Pan and Zeng, 2023, Rao et al., 2023]; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits [Barua et al., 2024, Huang et al., 2023d, Romero et al., 2024].
@@
Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024, Dorner et al., 2023, Gupta et al., 2024, Pellert et al., 2024, Shu et al., 2024, Song et al., 2023, Serapio-García et al., 2023], Some researchers have adapted them for more real-world scenarios. For example, Bhandari et al. [2025a], Frisch and Giulianelli [2024], Klinkert et al. [2024], Liu et al. [2024a], Zheng et al. [2025], Zou et al. [2024], Song et al. [2024b] contextualize the tests in topic-specific or open-domain conversations. Ai et al. [2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs. Meanwhile, Jiang et al. [2023] compile and adapt existing tests into the Machine Personality Inventory, while Mao et al. [2024a] develop the PersonalityEdit inventory using LLMs. However, Peereboom et al. [2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis (§ 7.2.2).
@@
The Big Five offers robust empirical foundations and is the most prominent model for personality assessment. However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021]. HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025]. In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales. While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005]. The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
Main Findings.
Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024]. Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data. More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a, Karra et al., 2022, La Cava and Tagarelli, 2024, Huang et al., 2023e, Klinkert et al., 2024]. The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024]. Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Zhang et al., 2024a, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities.
Personality differences between models are noteworthy. Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022, Jiang et al., 2023, La Cava and Tagarelli, 2024, Bodroža et al., 2024, Huang et al., 2023e, Klinkert et al., 2024]. LLM personalities also vary across contexts. The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023, Caron and Srivastava, 2022, Bodroža et al., 2024, Petrov et al., 2024]. This variability challenges the applicability of human personality theories that assume relative trait stability. Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022, Gupta et al., 2024, Dorner et al., 2023]. We defer related discussions to § 7.

### v2

@@ Abstract
The advancement of large language models (LLMs) has outpaced traditional evaluation methodologies. This progress presents novel challenges, such as measuring human-like psychological constructs, moving beyond static and task-specific benchmarks, and establishing human-centered evaluation. These challenges intersect with psychometrics, the science of quantifying the intangible aspects of human psychology, such as personality, values, and intelligence. This review paper introduces and synthesizes the emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs. The reviewed literature systematically shapes benchmarking principles, broadens evaluation scopes, refines methodologies, validates results, and advances LLM capabilities. Diverse perspectives are integrated to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field. Ultimately, the review provides actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit. A curated repository of LLM psychometric resources is available at https://github.com/valuebyte-ai/Awesome-LLM-Psychometrics.
@@ 1 Introduction
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024, Ye et al., 2024]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
Traditional AI evaluation has relied on curating task-specific datasets, annotating ground-truth labels with human input, running models on these datasets, and assessing performance using predefined metrics. However, LLMs have triggered an evaluation crisis, as their versatile capabilities and human-like behaviors exceed what traditional benchmarks can measure. Novel challenges include but are not limited to (1) evaluating psychological constructs like personality, values, and cognitive biases; (2) obsolescence of static benchmarks due to rapid LLM development and training data contamination; (3) compromised robustness and validity due to LLMs’ prompt- and context-sensitivity; (4) requiring human-centered evaluation approaches; and (5) expanding evaluation methodologies in scope and complexity as LLMs integrate into agentic and multimodal systems.
@@
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [Pasquali, 2009]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014]. The intersection of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
@@
This intersection sets the stage for a new research frontier. We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (e.g., personality, values, morality, and attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities). Research in LLM Psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs. Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor. By measuring and elucidating psychological constructs in LLMs, LLM Psychometrics further informs strategies for their targeted enhancement.
@@
The field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities. Despite this diversity, there is a lack of cohesion among researchers, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs. Consequently, there is a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated understanding of the field.
@@
Taxonomy and paper structure.
To bridge the gap, we systematically review LLM Psychometrics across evaluation, validation, and enhancement. The evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve). Accordingly, this review is structured as follows (Fig. 1). § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. § 3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks. The core measurement framework is detailed in § 4, § 5, and § 6. § 4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In § 5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 6, which examines the psychometric validation of evaluation results. Beyond evaluation, § 7 introduces strategies for enhancing LLMs through psychometric insights. § 8 discusses emerging trends, challenges, and future directions. Finally, § 9 highlights key ethical considerations, while § 10 concludes the paper.
Scope and related work.
We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024]. Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025]. Hagendorff [2023], Hagendorff et al. [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic review of LLM Psychometrics.
@@ 4 Psychometrics for measuring psychological constructs
HEXACO Wang et al. [2025b], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
@@
ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al. [2022]
@@ 4.1 Measuring personality constructs
LLMs exhibit personality constructs that are not explicitly programmed or trained towards. The emergent constructs critically shape LLM behavior [Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024]. Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
Table 2 and Table 3 summarize the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, Li et al. [2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
@@ 4.1.1 Personality traits
Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [APA Dictionary of Psychology, ]. In the context of LLMs, personality traits relate to model safety, bias, and toxicity [Zhang et al., 2024a, Wang et al., 2025b], and highly determine user experience [Klinkert et al., 2024].
@@
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [Goldberg, 2013], HEXACO [Ashton and Lee, 2007], Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], and Dark Triad [Paulhus and Williams, 2002]. When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits; MBTI assessment [Myers, 1985] for MBTI types; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits.
@@
Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024], others have adapted them for more real-world scenarios. For example, Bhandari et al. [2025a] contextualize the tests in topic-specific or open-domain conversations; Ai et al. [2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; Jiang et al. [2023] compile and adapt existing tests into the Machine Personality Inventory; and Mao et al. [2024a] develop the PersonalityEdit inventory using LLMs. However, Peereboom et al. [2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
Main findings.
Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024]. Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data. More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a]. The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024]. Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities. Personality differences between models are noteworthy. Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022]. LLM personalities also vary across contexts. The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023]. This variability challenges the applicability of human personality theories that assume relative trait stability. Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022].
@@
The Big Five offers robust empirical foundations and is the most prominent model for personality assessment. However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021]. HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025b]. In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales. While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005]. The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].